FIND THE TWO-HUNDREDTH DIGIT IN THE DECML REPRESENTATION OF EACH OF THE FOLLOWING FRACTIONS:
5/6
4/11
1/13
: FIND THE TWO-HUNDREDTH DIGIT IN THE DECML REPRESENTATION OF EACH OF THE FOLLOWING FRACTIONS:
5/6
4/11
1/13

Answer by checkley77(3848)

5/6=1.333333(3) answer=3.
4/11=.3636363(6) answer=6.
1/13=.0769230(7)6923 answer=7